OpenStudy (anonymous):

x and y intercepts of 2x-4y=12

OpenStudy (anonymous):

The y intercept is (0,-12) and the x intercept is (24,0)

OpenStudy (johnweldon1993):

You would make 'x' = 0 in your equation to find the y-intercept ...and you would make 'y' = 0 in your equation to find the x-intercept 2x - 4y = 12 x-intercept 2x - 4(0) = 12 2x = 12 x = 6 your x-intercept is (6,0) y-intercept 2x - 4y = 12 2(0) - 4y = 12 -4y = 12 y = -3 your y-intercepts is (0,-3)